A boy is 12 years old. His mother is four times as old as he is. In x years time the mother will be three times as old as the son. Form an equation in x and hence find x
Answer by jhunjiro(67)


age of boy = 12
age of mother = 4 times age of boy = 4*12 = 48
let x = years time
3(12+x) = 48+x
36+3x = 48+x
3x-x = 48-36
2x = 12
x = 6
In 6 years time the mother will be three times as old as the son.